please help me solve this question. a dog trainer has 108 ft of fencing that will be used to create a rectangular work area for dogs. If the trainer wants to enclose an area of 608ft what will be the dimensions of the work area
Answer by Alan3354(69443)

a dog trainer has 108 ft of fencing that will be used to create a rectangular work area for dogs. If the trainer wants to enclose an area of 608ft what will be the dimensions of the work area
------------
P = 2L + 2W = 108
L + W = 54
-------------
Area = L*W = 608 --- not ft, square feet
===
Find a pair of factors of 608 with a sum of 54
---
1*608 NG
2*304 NG
etc